# Alleviating Distribution Shift in Synthetic Data for Machine Translation Quality Estimation

### v2

@@ Abstract
Quality Estimation (QE) models evaluate the quality of machine translations without reference translations, serving as the reward models for the translation task. Due to the data scarcity, synthetic data generation has emerged as a promising solution. However, synthetic QE data often suffers from distribution shift, which can manifest as discrepancies between pseudo and real translations, or in pseudo labels that do not align with human preferences. To tackle this issue, we introduce ADSQE, a novel framework for alleviating distribution shift in synthetic QE data. To reduce the difference between pseudo and real translations, we employ the constrained beam search algorithm and enhance translation diversity through the use of distinct generation models. ADSQE uses references—i.e., translation supervision signals—to guide both the generation and annotation processes, enhancing the quality of token-level labels. ADSE further identifies the shortest phrase covering consecutive error tokens, mimicking human annotation behavior, to assign the final phrase-level labels. Specially, we underscore that the translation model can not annotate translations of itself accurately. Extensive experiments demonstrate that ADSQE outperforms SOTA baselines like CometKiwi in both supervised and unsupervised settings. Further analysis offers insights into synthetic data generation that could benefit reward models for other tasks.
@@ 1 Introduction
In this paper, we propose a framework, ADSQE (as shown in Figure 1), to Alleviate the Distribution Shift problem when generating synthetic QE data. Specifically, we first train two translation models as Generator and Annotator respectively. We then use the Generator to generate synthetic translations with constrained beam search (CBS) Geng et al. (2023b), which preserves the main structure of references while maximizing generation probabilities. This allows us to treat matched tokens as correct with high accuracy using TER Snover et al. (2006) tool. For the mismatched part, we utilize generation probabilities provided by the Annotator to rejudge their fine-grained severities. Since human annotators typically prefer annotating entire phrases as spans, we propose an algorithm, which helps aggregate token-level labels into phrase-level labels.
Extensive experiments across three language directions (English-German, Chinese-English, and Hebrew-English) demonstrate that ADSQE achieves new (SOTA) results in both supervised and unsupervised settings. Further analysis offers insights into synthetic data generation that may benefit general reward models: (1) distribution shift problem is crucial in synthetic data methods; (2) diversity between annotation and generation models enhances annotation accuracy; (3) diversity in generation models provides further improvements; (4) the capacity of the generation model must be balanced; (5) enhancing the capacity of the annotation model with supervision signals is helpful.
To promote further research in this field, we will provide open access to our code and datasets (please refer to supplementary materials).
@@ 3.2 Synthetic Labels
Leveraging supervision signals. When the model lacks the necessary translation knowledge for the given input, its generation probabilities become speculative and fail to meaningfully correlate with the severity of errors. To address this issue, we ensure that the translation model is trained on parallel pairs that are also used for generating synthetic data. Therefore, the Annotator becomes “professional” when working with the given parallel pairs, rather than “amateur”.
@@
First, we parse the synthetic translation to obtain its dependency tree. Consecutive “BAD” tokens are added to the candidate set, which serves as the initial phrase (i.e. “action with his”). Following this, we iteratively execute the following steps:
@@
  • Step b. To ensure syntactic coherence, we add all tokens along the path from the LCA into the candidate set (i.e. the word “consent” is added as it lies on the path from “his” to “ take”).
  • Step c. To ensure the candidate tokens are consecutive, we add all tokens, which are located between the leftmost and rightmost candidate tokens, into the candidate set (i.e. the word “some” is added as it appears within the phrase “take some action with his consent”).
@@ 4 Experiment
In the experiment section, we aim to address the following research questions regarding ADSQE: (1) How does ADSQE perform across supervised and unsupervised settings for various language pairs? (2) To align synthetic labels with human preferences, we introduce the Annotator and the SPCE algorithm to generate MQM data. To what extent are these techniques effective? (3) Can the Annotator and the Generator be the same model? In other words, can a model fairly annotate its own output? (4) Does increasing the diversity of Generators lead to enhanced QE performance? (5) How does the translation performance of the Generator and the Annotator influence the quality of the synthetic data? (6) Does ADSQE demonstrate advantages in terms of generation cost and convergence efficiency?
To address Q1, we evaluate ADSQE in both supervised and unsupervised settings across multiple language pairs (EN-DE, ZH-EN, HE-EN). For Q2, we conduct ablation studies to quantify the individual contributions of the Annotator and the SPCE algorithm in improving the alignment of synthetic labels with human preferences. Regarding Q3, we investigate the feasibility of employing a single model to simultaneously serve as both the Generator and the Annotator. To explore Q4, we enhance the diversity of Generators by training them on distinct parallel subsets. To address Q5, we control the translation performance of the Generator and the Annotator by training them on corpora of varying sizes. Finally, for Q6, we compare ADSQE against other synthetic data approaches in terms of generation cost and convergence efficiency.
@@ 4.1 Experiment Setup
Setting Lang Model Name Sentence-level Word-level Span-level
@@
ADSQE 26.27 44.04 10.27 11.57 - - -
Table 2: Main results on different QE test sets. We follow the setting in GEMBA-MQM Kocmi and Federmann (2023), which utilizes a few-shot prompt containing examples for EN-DE and ZH-EN language pairs. However, since the prompt does not include examples for HE-EN, GEMBA-MQM is treated as a supervised method for EN-DE and ZH-EN, while an unsupervised one for HE-EN.
@@
Since CometKiwi does not provide sentence-level results for the individual model, we reproduce sentence-level results using the released model [^5] https://huggingface.co/Unbabel/wmt23-cometkiwi-da-xl. For word- and span-level results, we directly utilize the results provided in the CometKiwi report Rei et al. (2023), as the corresponding word- and span-level models have not been released. InstructScore is implemented using 10K synthetic data released by Xu et al. (2023). For ADSQE and MQMQE, we generate 500K synthetic data in each language pair. In Section 4.5, we demonstrate that ADSQE remains competitive with InstructScore when trained on 10K synthetic data while requiring significantly lower costs. Additional implementation details are provided in Appendix A.
@@ 4.2 Main Results
As demonstrated in Table 2, ADSQE substantially outperforms CometKiwi despite utilizing fewer parameters, achieving notable improvements with average gains of 4.38 in Spearman, 3.41 in MCC, and 1.45 in F1 score. Moreover, ADSQE significantly outperforms GEMBA-MQM, which is based on the advanced LLM, GPT-4. Compared to other synthetic methods, i.e., MQMQE and InstructScore, ADSQE demonstrates consistent superiority, indicating that our synthetic data is of higher quality.
@@
As shown in Table 2, both MQMQE and InstructScore demonstrate significant performance declines compared to their supervised counterparts, with reductions of 15.74 and 7.64 on average, respectively. That implies the distribution shift problem in previous synthetic QE data. In contrast, our proposed method achieves superior robustness, incurring a smaller average reduction of 6.64 points. Moreover, ADSQE outperforms CometKiwi, which relies on labeled datasets from other language pairs, on HE-EN. This indicates that our synthetic data provides more relevant QE knowledge for HE-EN compared to datasets from different language pairs.
@@ 4.3 Ablation Study
The ADSQE framework introduces two techniques to align the synthetic labels with human preferences: (1) leveraging the Annotator to rejudge tokens initially labeled as BAD by the TER tool, and (2) applying the SPCE Algorithm to aggregate token-level annotations into span-level annotations. The effectiveness of these techniques within the ADSQE framework is empirically validated in Table 3. The results demonstrate that the Annotator effectively corrects errors in coarse-grained labels. Furthermore, the SPCE algorithm successfully identifies phrase spans, thereby achieving better alignment with human preferences.
@@ 4.4 Analysis
In this subsection, we aim to investigate the contributions of the Generator and the Annotator to the performance of ADSQE. We create translation models with varying levels of performance by training them on datasets of different sizes. We train three translation models—$S$, $M$, and $L$—using 1M, 5M, and 20M distinct sentence pairs, respectively. To enhance diversity while maintaining similar performance, we then train three additional models—$S^{\prime}$, $M^{\prime}$, and $L^{\prime}$—on the remaining 1M, 5M, and 20M sentence pairs. All subsequent analyses are conducted on the WMT23 EN-DE test set in the unsupervised setting.
@@
Table 4: Analysis comparing individual and collaborative results of Generator $M$ and Annotator $L^{\prime}$.
The model cannot fairly annotate its outputs. As discussed in Section 3.2, the translation model may exhibit overconfidence in its output, leading to an increased number of “OK” labels. To investigate that, we calculate the error rate across different settings. Table 4 shows that translation models tend to consistently assign more “OK” labels to their own output, regardless of the model’s performance. The poor QE performance implies that most of the “OK” labels are false-negative.
@@
The capacity of the Generators must be balanced. If the Generator is too strong, the generated data will have few errors, limiting learning opportunities; if the Generator is too weak, the data will be too noisy or unrealistic. To investigate this point, we measure the error rate and similarity between synthetic translation and the translation across different Generators in Table 6. To be specific, the similarity is measured by the BLEU score, which calculates between synthetic translation and the “real” translation from the WMT2023 QE validation set.
@@
To examine the influence of error rate, we fixes Generator $M$ for generation and downsamples its synthetic data to match the error rates corresponding to each Generator. Likewise, to examine the influence of similarity, we fix the error rate of Generator $M$ for downsampling and utilize different Generators for generation. As shown in Figure 3, when the similarity is fixed, the performance degradation as the error rate decreases. Similarly, when the error rate is fixed, the performance improves with increasing similarity. That explains why ADSQE achieves the highest performance using Generator $M$.
Figure 3: Impact of Generator metrics on synthetic data quality in sentence-level and word-level tasks.
@@ 4.5 Generation Cost
We measure the generation cost for various synthetic data approaches using 10K samples. As shown in Table 8, ADSQE generates synthetic data 14.29$\times$ faster than InstructScore. Meanwhile, with the same amount of synthetic data, ADSQE also achieves a notable improvement of 14.29 in the Spearman score with slight decline of 1.76 in MCC. Despite introducing rejudge and SPCE algorithm, ADSQE does not substantially increase computational complexity, requiring only 3.12$\times$ the timelapse of MQMQE. Further analysis of ADSQE time overhead is provided in Section D.
@@
Table 8: Average generation time per sample of different pseudo data methods on a single V100 GPU.
@@ 6 Conclusion
The distribution shift of synthetic data poses a persistent challenge in the QE field. To tackle this challenge, we introduce the ADSQE framework, which mitigates distribution shifts by utilizing translation references (a form of supervision signal) and enhancing the diversity among translation models. Experiments show that ADSQE achieves SOTA results in both supervised and unsupervised settings. Furthermore, our analysis underscores some insights for synthetic data generation, which could benefit synthetic data methods for general reward models.
@@ A.2 Generator & Annotator
The model parameters are optimized using the Adam optimizer, configured with $\beta_{1}=0.9$ and $\beta_{2}=0.98$. The initial learning rate is set to 5e-4, and the inverse square root learning rate scheduler is employed with 6000 warm-up steps. Model parameters update every 20 batches and the batch is configured to 13,650. The dropout rate is set to 0.3, and the weight decay is set to 1e-4. The training objective employs the label-smoothed cross-entropy criterion with $eps=0.1$. The training stops early if there are no improvements in validation performance for the last 15 epochs.
@@ A.4 GEMBA-MQM
We adopt the same configuration as GEMBA-MQM Kocmi and Federmann (2023), but utilize the gpt-4-turbo-preview model, and structured QE data is extracted from the regular expression.
@@ Appendix B Details of SPCE algorithm
We utilize the Stanza toolkit Qi et al. (2020) to perform dependency tree parsing on the synthetic translations. The pseudo-code for the implementation of the SPCE algorithm is presented in Algorithm 1.
@@ Appendix C Case Study
Appendix D Time Analysis
ADSQE can be divided into four stages: (1) Synthesize Translations with the Generator, (2) Synthesize coarse-grained labels with TER tool, (3) Refining labels with the Annotator, (4) Aggregate the results with the SPCE algorithm. We counted the time spent on each stage in Table 11.
@@ Appendix D Time Analysis
Table 11: Time Analysis of ADSQE.

### v3

@@ Abstract
Quality Estimation (QE) models evaluate the quality of machine translations without reference translations, serving as the reward models for the translation task. Due to the data scarcity, synthetic data generation has emerged as a promising solution. However, synthetic QE data often suffers from distribution shift, which can manifest as discrepancies between pseudo and real translations, or in pseudo labels that do not align with human preferences. To tackle this issue, we introduce DCSQE, a novel framework for alleviating distribution shift in synthetic QE data. To reduce the difference between pseudo and real translations, we employ the constrained beam search algorithm and enhance translation diversity through the use of distinct generation models. DCSQE uses references—i.e., translation supervision signals—to guide both the generation and annotation processes, enhancing the quality of token-level labels. DCSQE further identifies the shortest phrase covering consecutive error tokens, mimicking human annotation behavior, to assign the final phrase-level labels. Specially, we underscore that the translation model can not annotate translations of itself accurately. Extensive experiments demonstrate that DCSQE outperforms SOTA baselines like CometKiwi in both supervised and unsupervised settings. Further analysis offers insights into synthetic data generation that could benefit reward models for other tasks. The code is available at https://github.com/NJUNLP/njuqe.
@@ 1 Introduction
In this paper, we propose a framework, DCSQE (as shown in Figure 1), Distribution-Controlled Data Synthesis for QE, aiming to alleviate the distribution shift problem when generating synthetic QE data. Specifically, we first train two translation models as Generator and Annotator respectively. We then use the Generator to generate synthetic translations with constrained beam search (CBS) Geng et al. (2023b), which preserves the main structure of references while maximizing generation probabilities. This allows us to treat matched tokens as correct with high accuracy using TER Snover et al. (2006) tool. For the mismatched part, we utilize generation probabilities provided by the Annotator to rejudge their fine-grained severities. Since human annotators typically prefer annotating entire phrases as spans, we propose an algorithm, which helps aggregate token-level labels into phrase-level labels.
Extensive experiments across three language directions (English-German, Chinese-English, and Hebrew-English) demonstrate that DCSQE achieves new (SOTA) results in both supervised and unsupervised settings. Further analysis offers insights into synthetic data generation that may benefit general reward models: (1) distribution shift problem is crucial in synthetic data methods; (2) diversity between annotation and generation models enhances annotation accuracy; (3) diversity in generation models provides further improvements; (4) the capacity of the generation model must be balanced; (5) enhancing the capacity of the annotation model with supervision signals is helpful.
@@ 3.2 Synthetic Labels
Leveraging supervision signals. When the model lacks the necessary translation knowledge for the given input, its generation probabilities become speculative and fail to meaningfully correlate with the severity of errors. To address this issue, we ensure that the Annotator is trained on parallel pairs that are also used for generating synthetic data. Therefore, the Annotator becomes “professional” when working with the given parallel pairs, rather than “amateur”.
@@
First, we parse the synthetic translation to obtain its dependency tree. Consecutive “BAD” tokens are added to the candidate set, which serves as the initial phrase. Following this, we iteratively execute the following steps:
@@
  • Step b. To ensure syntactic coherence, we add all tokens along the path from the LCA into the candidate set (i.e. word “consent” which is on the path from “his” to “take”).
  • Step c. To ensure the candidate tokens are consecutive, we add all tokens, which are located between the leftmost and rightmost candidate tokens, into the candidate set (i.e. word “some” which is in the phrase “take some action with his consent”).
@@ 4 Experiment
In the experiment section, we aim to address the following research questions regarding DCSQE: (1) How does DCSQE perform across supervised and unsupervised settings for various language pairs? (2) To align synthetic labels with human preferences, we introduce the Annotator and the SPCE algorithm to generate MQM data. To what extent are these techniques effective? (3) Can the Annotator and the Generator be the same model? In other words, can a model fairly annotate its own output? (4) Does increasing the diversity of Generators lead to enhanced QE performance? (5) How does the translation performance of the Generator and the Annotator influence the quality of the synthetic data? (6) Does DCSQE demonstrate advantages in terms of generation cost and convergence efficiency?
To address Q1, we evaluate DCSQE in both supervised and unsupervised settings across multiple language pairs (EN-DE, ZH-EN, HE-EN). For Q2, we conduct ablation studies to quantify the individual contributions of the Annotator and the SPCE algorithm in improving the alignment of synthetic labels with human preferences. Regarding Q3, we investigate the feasibility of employing a single model to simultaneously serve as both the Generator and the Annotator. To explore Q4, we enhance the diversity of Generators by training them on distinct parallel subsets. To address Q5, we control the translation performance of the Generator and the Annotator by training them on corpora of varying sizes. Finally, for Q6, we compare DCSQE against other synthetic data approaches in terms of generation cost.
@@ 4.1 Experiment Setup
Setting Dataset Method Sentence-level Word-level Span-level
@@
DCSQE 26.27 44.04 10.27 11.57 - - -
Table 2: Main results on different QE test sets. We follow the setting in GEMBA-MQM Kocmi and Federmann (2023), which utilizes a few-shot prompt containing examples for EN-DE and ZH-EN language pairs. However, since the prompt does not include examples for HE-EN, GEMBA-MQM is treated as a supervised method for EN-DE and ZH-EN, while an unsupervised one for HE-EN. The details of CometKiwi are described in the second paragraph of Section 4.1.
@@
Since CometKiwi does not provide sentence-level results for the individual model, we reproduce sentence-level results using the released CometKiwi-XL [^5] https://huggingface.co/Unbabel/wmt23-CometKiwi-da-xl. For word- and span-level results, we directly utilize the results provided in the CometKiwi report Rei et al. (2023) (where model size is greater than or equal to XL), as the corresponding word- and span-level models have not been released. InstructScore is implemented using 10K synthetic data released by Xu et al. (2023). For DCSQE and MQMQE, we generate 500K synthetic data in each language pair. In Section 4.5, we demonstrate that DCSQE remains competitive with InstructScore when trained on 10K synthetic data while requiring significantly lower costs. Additional implementation details are provided in Appendix A.
@@ 4.2 Main Results
As demonstrated in Table 2, DCSQE substantially outperforms CometKiwi despite utilizing fewer parameters, achieving notable improvements with average gains of 4.38 in Spearman, 3.41 in MCC, and 1.45 in F1 score. Moreover, DCSQE significantly outperforms GEMBA-MQM, which is based on the advanced LLM, GPT-4. Compared to other synthetic methods, i.e., MQMQE and InstructScore, DCSQE demonstrates consistent superiority, indicating that our synthetic data is of higher quality.
@@
As shown in Table 2, both MQMQE and InstructScore demonstrate significant performance declines compared to their supervised counterparts, with reductions of 15.74 and 7.64 on average, respectively. That implies the distribution shift problem in previous synthetic QE data. In contrast, our proposed method achieves superior robustness, incurring a smaller average reduction of 6.64 points. Moreover, DCSQE outperforms CometKiwi, which relies on labeled datasets from other language pairs, on HE-EN. This indicates that our synthetic data provides more relevant QE knowledge for HE-EN compared to datasets from different language pairs.
@@ 4.3 Ablation Study
The DCSQE framework introduces two techniques to align the synthetic labels with human preferences: (1) leveraging the Annotator to rejudge tokens initially labeled as BAD by the TER tool, and (2) applying the SPCE Algorithm to aggregate token-level annotations into span-level annotations. The effectiveness of these techniques within the DCSQE framework is empirically validated in Table 3. The results demonstrate that the Annotator effectively corrects errors in coarse-grained labels. Furthermore, the SPCE algorithm successfully identifies phrase spans, thereby achieving better alignment with human preferences.
@@ 4.4 Analysis
In this subsection, we aim to investigate the contributions of the Generator and the Annotator to the performance of DCSQE. We create translation models with varying levels of performance by training them on datasets of different sizes. We train three translation models—$S$, $M$, and $L$—using 1M, 5M, and 20M distinct sentence pairs, respectively. To enhance diversity while maintaining similar performance, we then train three additional models—$S^{\prime}$, $M^{\prime}$, and $L^{\prime}$—on the remaining 1M, 5M, and 20M sentence pairs. All subsequent analyses are conducted on the WMT23 EN-DE test set in the unsupervised setting.
@@
Table 4: Analysis comparing individual and collaborative results of deploying models $L$ and $M$.
The model cannot fairly annotate its outputs. As discussed in Section 3.2, the translation model may exhibit overconfidence in its output, leading to an increased number of “OK” labels. To investigate that, we calculate the error rate across different settings. Table 4 shows that translation models tend to consistently assign more “OK” labels to their own output, regardless of the model’s performance. The poor QE performance implies that most of the “OK” labels are false-negative. We also provide a corresponding case study (Table 10 see in Appendix).
@@
The capacity of the Generators must be balanced. If the Generator is too strong, the generated data will have few errors, limiting learning opportunities; if the Generator is too weak, the data will be too noisy or unrealistic. To investigate this point, we measure the Error Rate of synthetic translations and the Similarity between synthetic and real translations across different Generators in Table 6. To be specific, the Similarity is measured by the BLEU score, which calculates between synthetic translations and the “real” translations from the WMT2023 QE validation set.
@@
Figure 3: Impact of Generator metrics on synthetic data quality in sentence-level and word-level tasks. The first row represents the sentence-level results. The second row represents the word-level results.
To isolate the individual impact of each factor on QE performance , we employ the following downsampling strategy to select subsets of translations from one Generator—referred to as the original Generator—that match the Error Rate distribution of another Generator. Specifically, we begin by sampling a set of Error Rate scores from the Error Rate distribution of one Generator. For each sampled Error Rate score, we select a translation, whose Error Rate score is closest to the sampled score, from the original Generator. All samples are drawn without replacement. This downsampled subset retains the translation style (Similarity) of the original Generator while mirroring the Error Rate distribution of another Generator.
We consider three experimental settings. In the “Fixed Similarity”, we fix the Similarity by exclusively deploying model M as the original Generator and downsample three subsets to mirror the Error Rate distribution of S, M, and L. Conversely, in the “Fixed Error Rate”, we exclusively mirror the Error Rate distribution of M and downsample data from original Generators S, M, and L to diversify the Similarity level. We also include the “DCSQE” for comparison. As shown in Figure 3, when the Similarity is fixed, the performance degradation as the Error Rate decreases. Similarly, when the Error Rate is fixed, the performance improves with increasing Similarity. That explains why DCSQE achieves the highest performance using Generator $M$.
@@ 4.5 Generation Cost
We measure the generation cost for various synthetic data approaches using 10K samples. As shown in Table 8, DCSQE generates synthetic data 14.29$\times$ faster than InstructScore. Meanwhile, with the same amount of synthetic data, DCSQE also achieves a notable improvement of 14.29 in the Spearman score with slight decline of 1.76 in MCC. Despite introducing rejudge and SPCE algorithm, DCSQE does not substantially increase computational complexity, requiring only 3.12$\times$ the timelapse of MQMQE. Further analysis of DCSQE time overhead is provided in Section D.
@@
Table 8: Average generation time per sample of different synthetic data methods on a single V100 GPU.
@@ 6 Conclusion
The distribution shift of synthetic data poses a persistent challenge in the QE field. To address this challenge, we propose the DCSQE framework, which mitigates distribution shift by leveraging translation references, a form of translation supervision signal, to guide the generation of both diverse synthetic translations and their corresponding synthetic labels. Experiments show that DCSQE achieves SOTA results in both supervised and unsupervised settings. Furthermore, our analysis underscores some insights for synthetic data generation, which could benefit synthetic data methods for general reward models.
@@
Acknowledgments
We would like to thank the anonymous reviewers for their insightful comments. Shujian Huang is the corresponding author. This work is supported by National Science Foundation of China (No. 62376116, 62176120), the Fundamental Research Funds for the Central Universities (No. 2024300507).
@@ A.2 Generator & Annotator
The model parameters are optimized using the Adam optimizer, configured with $\beta_{1}=0.9$ and $\beta_{2}=0.98$. The initial learning rate is set to 5e-4, and the inverse square root learning rate scheduler is employed with 6000 warm-up steps. Model parameters update every 20 batches and the batch is configured to 13,650. The dropout rate is set to 0.3, and the weight decay is set to 1e-4. The training objective employs the label-smoothed cross-entropy criterion with $\text{eps}=0.1$. The training stops early if there are no improvements in validation performance for the last 15 epochs.
@@ A.4 GEMBA-MQM
We adopt the same configuration as GEMBA-MQM Kocmi and Federmann (2023), but utilize the gpt-4-0125-preview model, and structured QE data is extracted from the regular expression.
@@ Appendix B Details of SPCE algorithm
The pseudo-code for the implementation of the SPCE algorithm is presented in Algorithm 1.
The SPCE algorithm depends on models for dependency parsing. However, compared to QE tasks, dependency parsing is a well-established NLP task that has been extensively studied for decades, with significantly more available resources. For instance, the Universal Dependencies (UD) project Nivre et al. (2020) covers over 150 languages with 200 treebanks. To further imporve the performance on low-resource languages, TowerParse Glavaš and Vulić (2021) leverages a heuristic data selection strategy to automatically identify suitable training data from UD for any target language, achieving competitive results. In this work, we utilize the Stanza toolkit Qi et al. (2020) to perform dependency parsing on the synthetic translations.
@@ Appendix C Case Study
Section 4 demonstrates that the model exhibits overconfidence in its own outputs, resulting in erroneous “OK” labels. For example, in Table 10, the words “sind” and “einem” are incorrectly labeled as “OK” by model $M$ itself, despite being grammatical errors. Deploying another model $M^{\prime}$ as the Annotator helps mitigate this issue. Leveraging the SPCE algorithm, the phrase “sind in einem” is correctly identified as a complete unit and labeled as “BAD”. "privaten Vorstellung" and “Privatshow” are synonyms, and the former has been correctly rejudged as “OK”.
Source Helena Anderson is in a private show
Reference Helena Anderson sich in einer Privatshow
Generator $M$ Helena Anderson sind in einem privaten Vorstellung
Annotator $M$ OK OK OK OK OK OK OK
Annotator $M^{\prime}$ OK OK BAD BAD BAD OK OK
Table 10: A Case Study illustrating the model exhibits overconfidence in its own outputs. Model $M$ consistently assigning the OK labels to all tokens generated by itself.
Section 3 demonstrates that supervision signals is helpful for enhancing the capacity of annotation model. In Table 11, the word “Anträge” is correctly labeled as “BAD” by model $M_{\text{amateur}}$. However, the assigned error severity is incorrect. Since Annotator $M^{\prime}$ has been trained on this specific parallel sentence, it provides the correct error severity and additionally identifies another error, the word “aufgenommen”. Furthermore, Annotator $L^{\prime}$, possessing more advanced translation capability, produces entirely accurate annotations.
@@
Table 12: Time Analysis of DCSQE.
Translation Error Type w/o Synthetic Data w/ Synthetic Data
accuracy/addition 74.22% 85.39%
accuracy/creative reinterpretation 21.42% 35.57%
accuracy/mistranslation 43.50% 59.23%
accuracy/source language fragment 46.09% 55.87%
fluency/grammar 31.31% 47.79%
fluency/inconsistency 26.60% 38.38%
fluency/punctuation 32.64% 42.39%
fluency/register 27.00% 39.00%
fluency/spelling 32.91% 43.18%
locale convention/currency format 56.67% 56.67%
locale convention/date format 50.00% 100.00%
locale convention/time format   0.00% 40.00%
style/bad sentence structure 31.95% 38.96%
style/unnatural or awkward 32.01% 43.47%
terminology/inappropriate for context 27.74% 38.41%
terminology/inconsistent 26.67% 33.33%
non-translation! 26.38% 63.83%
other 43.07% 53.92%
Table 13: Comparison of error detection accuracy with and without synthetic data across translation error types.
Appendix D Time Analysis
DCSQE can be divided into four stages: (1) Synthesize Translations with the Generator, (2) Synthesize coarse-grained labels with TER tool, (3) Refining labels with the Annotator, (4) Aggregate the results with the SPCE algorithm. We counted the time spent on each stage in Table 12.